Question 6

Drake Corporation is reviewing an investment proposal. The initial cost is $103,100. Estimates of the book value of the investment at the end of each year, the net cash flows for each year, and the net income for each year are presented in the schedule below. All cash flows are assumed to take place at the end of the year. The salvage value of the investment at the end of each year is assumed to equal its book value. There would be no salvage value at the end of the investment’s life.

Investment Proposal
Year Book Value Annual
Cash Flows
Annual
Net Income
1 $69,700 $46,000 $12,600
2 41,100 40,100 11,500
3 20,900 34,000 13,800
4 7,000 29,200 15,300
5 0 25,970 18,970


Drake Corporation uses an 11% target rate of return for new investment proposals.

Click here to view PV table.

(a)

What is the cash payback period for this proposal? (Round answer to 2 decimal places, e.g. 10.50.)

Cash payback period

years


(b)

What is the annual rate of return for the investment? (Round answer to 2 decimal places, e.g. 10.50%.)

Annual rate of return for the investment

%


(c)

What is the net present value of the investment? (If the net present value is negative, use either a negative sign preceding the number e.g. -45 or parentheses eg (45). Round answer to 0 decimal places, e.g. 125. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Net present value $

(c)

What is the net present value of the investment? (If the net present value is negative, use either a negative sign preceding the number e.g. -45 or parentheses eg (45). Round answer to 0 decimal places, e.g. 125. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Net present value $
Answer #1

Solution a:

Computation of cumulative cash flows
Year Cash Flows Cumulative cash flows
1 $46,000.00 $46,000.00
2 $40,100.00 $86,100.00
3 $34,000.00 $1,20,100.00
4 $29,200.00 $1,49,300.00
5 $25,970.00 $1,75,270.00

Payback period = 2 year + ($103100- 86100)/ $34000 = 2.50 years

Solution b:

Computation of NPV - Drake Company
Particulars Amount Period PV Factor Present Value
Cash Outflows:
Initial investment $1,03,100.00 0 1 $1,03,100
Present Value of Cash Outflows (A) $1,03,100
Cash Inflows:
Year 1 $46,000.00 1 0.90090 $41,441
Year 2 $40,100.00 2 0.81162 $32,546
Year 3 $34,000.00 3 0.73119 $24,860
Year 4 $29,200.00 4 0.65873 $19,235
Year 5 $25,970.00 5 0.59345 $15,412
Present Value of Cash Inflows (B) $1,33,495
Net Present Value (B-A) $30,395

Solution c:

Annual rate of return = Average annual income / Average investment

Average annual income = ($12,600 + $11,500 + $13,800 + $15,300 + $18,970)/5 = $14,434

Average investment = (Initial investment  + Salvage value)/2 = ($103,100 + 0) /2 = $51,550

Annual rate of return = 14434 / 51550 = 28.00%